# Extreme events and power-law distributions from nonlinear quantum dissipation

### v1

@@ II.1 Moments of the number distribution
Let us consider the case of $M=2$, which is the minimal model for observing the power-law behavior. We denote the excitation number operator as $\hat{n}=\hat{a}^{{\dagger}}\hat{a}$. Setting $\hat{O}=\hat{a}^{{\dagger}}\hat{a}$ in Eq. (12), and using the commutation relation $[\hat{a},\hat{a}^{{\dagger}}]=1$, yields the equation of motion
$\frac{d\braket{\hat{n}}}{dt}=\kappa_{1}+4\kappa_{2}+(-\gamma_{1}+\kappa_{2}+2\gamma_{2}+6\kappa_{2})\braket{\hat{n}}-2(\gamma_{2}-\kappa_{2})\braket{\hat{n}^{2}}.$ (13)
At the critical point, i.e, $\gamma_{2}=\kappa_{2}$, the mean excitation number $\braket{\hat{n}(t)}$ can be solved exactly to give
@@
Using the identities $[\hat{a},\hat{a}^{{\dagger}k}]=k\hat{a}^{{\dagger}k-1}$ and $[\hat{a}^{2},\hat{a}^{{\dagger}k}]=2k\hat{a}^{{\dagger}k-1}\hat{a}+k(k-1)\hat{a}^{{\dagger}k-2}$, which can be derived from the bosonic commutation relation, we obtain
@@
II.2 Higher-order bosonic correlation function
@@
We can derive general asymptotic expressions for $\braket{\hat{a}^{{\dagger}k}\hat{a}^{k}}$. Observe that for even $k$, the dominant contribution on the right hand side of Eq. (22) comes from $\braket{\hat{a}^{{\dagger}k-2}\hat{a}^{k-2}}$. Therefore, setting $k-2l$ we have the asymptotic recurrence relation for the even moments:
@@
for large $\nu$. This can be used potentially as an experimental signature of the power-law behavior.
II.3 Wigner quasiprobability distribution
@@
satisfying the canonical commutation relation $[\hat{q},\hat{p}]=i$. If the system physically describes a quantum particle trapped in a one-dimensional potential, $\hat{q}$ and $\hat{p}$ are the dimensionless position and momentum operators.
@@
$\begin{split}&\hat{a}\rho\longrightarrow\left(\alpha+\frac{1}{2}\partial_{\alpha^{*}}\right)W(\alpha,\alpha^{*})\equiv D_{1}W(\alpha,\alpha^{*}),\quad\rho\hat{a}^{{\dagger}}\longrightarrow D_{1}^{*}W(\alpha,\alpha^{*})\\ &\rho\hat{a}\longrightarrow\left(\alpha-\frac{1}{2}\partial_{\alpha^{*}}\right)W(\alpha,\alpha^{*})\equiv D_{2}W(\alpha,\alpha^{*}),\quad\hat{a}^{{\dagger}}\rho\longrightarrow D_{2}^{*}W(\alpha,\alpha)\end{split}$ (31)
@@
$0=(\gamma_{1}-\kappa_{1})(1+R\partial_{R})W+\frac{\gamma_{1}+\kappa_{1}}{2}\partial_{R}R\partial_{R}W+(\gamma_{2}-\kappa_{2})\left(4R+2R^{2}\partial_{R}+\frac{1}{2}R\partial_{R}^{2}R\partial_{R}\right)W+2(\gamma_{2}+\kappa_{2})(R\partial_{R}+R\partial_{R}R\partial_{R})W.$ (37)
@@
III Scaling behavior of the number statistics
Here, we derive analytical predictions for the divergences of the various statistics presented in the main text, such as the mean number of excitation quanta and the second-order correlation function, as we approach the critical point $\delta\to 0^{+}$. While the numerical results shown in the main text are for special case of $M=2$, the analytical treatment here holds for the general $M$-boson model.
@@ III Scaling behavior of the number statistics
which vanishes as $\delta^{1-\nu}$ for $10<\nu<$ 1 and is finite for $\nu>1$. For $k\in\mathbb{Z}^{+}$, the $k$-th moment of the distribution $p(n)$ is given by
[…]
@@
$g^{(2)}(0)=\frac{\braket{n^{2}}-\braket{n}}{\braket{n^{2}}}\sim\begin{cases}\text{finite},\quad&0<\nu<1\\[8.61108pt] \log(1/\delta),\quad&\nu=1\\[8.61108pt] \dfrac{1}{\delta^{\nu-1}},\quad&1<\nu<2\\[8.61108pt] \dfrac{1}{\delta^{3-\nu}\log^{2}(1/\delta)},\quad&\nu=2\\[8.61108pt] \log(1/\delta),\quad&\nu=3\\[8.61108pt] \text{finite},\quad&\nu>3\end{cases}$ (48)
which is divergent in the regime $1\leq\nu\leq 3$. This is consistent with the numerical results shown in the main text.
IV Quantum Liénard system
The quantum Liénard system, proposed in Refs. [21, 22], serves as a quantum mechanical analog to the (classical) Liénard system, which is well-studied in the context of nonlinear dynamics. The general Liénard system is given by the dynamical equations
@@ IV Quantum Liénard system
[…]
@@
IV.1 ‘Quantization’ of dissipative dynamical systems
Eq. (50) cannot be simulated directly, since the system of equations is not closed in general. Instead, we seek a generator of infinitesimal time translation in the form of a Lindblad master equation. A general construction of the master equation (referred to as ‘quantization’) for an arbitrary polynomial $f$ and $g$ can be found in Refs. [21, 22]. For completeness, we briefly summarize the steps involved in the ‘quantization’ procedure:
@@
    Since $\alpha$ is a classical variable (sometimes called a $c$-number), we can change the order of $\alpha$ and $\alpha^{*}$ freely in $f_{1}(\alpha,\alpha^{*})$ and $f_{2}(\alpha,\alpha^{*})$. For convenience, reorder the terms such that all the $\alpha^{*}$ appears on the left of $\alpha$.
@@
    The normal ordering of operators comes from Step 3. The choice of ordering here is simply for convenience. In principle, the master equation can be constructed for any arbitrary operator ordering.
@@
    for a suitable (and in general non-unique) choice of Hamiltonian $\hat{H}$, jump operators $\hat{c}_{j}$ and decay rates $\Gamma_{j}>0$, such that Eq. (54) is satisfied. The solution to this inverse problem is the subject of Refs. [21, 22], which also proved the existence of a master equation for any arbitrary polynomial $f_{1}$ and $f_{2}$.
IV.2 Example: Cubic Liénard system
The application of the ‘quantization’ procedure to the cubic Liénard system (51) is detailed in Refs. [21, 22]. The result is the Lindblad master equation
@@ IV.2 Example: Cubic Liénard system
[…]

### v2

@@ B.1 Moments of the number distribution
Let us consider the case of $M=2$, which is the minimal model for observing the power-law behavior. We denote the excitation number operator as $\hat{n}=\hat{a}^{{\dagger}}\hat{a}$. Setting $\hat{O}=\hat{a}^{{\dagger}}\hat{a}$ in Eq. (29), and using the commutation relation $\hat{a},\hat{a}^{{\dagger}}]=1$, yields the equation of motion
$\frac{d\braket{\hat{n}}}{dt}=\kappa_{1}+4\kappa_{2}+(-\gamma_{1}+\kappa_{1}+2\gamma_{2}+6\kappa_{2})\braket{\hat{n}}-2(\gamma_{2}-\kappa_{2})\braket{\hat{n}^{2}}.$ (30)
At the critical point, i.e., $\gamma_{2}=\kappa_{2}$, the mean excitation number $\braket{\hat{n}(t)}$ can be solved exactly to give
@@
Using the identities $\hat{a},\hat{a}^{{\dagger}k}]=k\hat{a}^{{\dagger}k-1}$ and $\hat{a}^{2},\hat{a}^{{\dagger}k}]=2k\hat{a}^{{\dagger}k-1}\hat{a}+k(k-1)\hat{a}^{{\dagger}k-2}$, which can be derived from the bosonic commutation relation, we obtain
@@
B.2 Higher-order bosonic correlation function
@@
We can derive general asymptotic expressions for $\braket{\hat{a}^{{\dagger}k}\hat{a}^{k}}$. Observe that for even $k$, the dominant contribution on the right hand side of Eq. (39) comes from $\braket{\hat{a}^{{\dagger}k-2}\hat{a}^{k-2}}$. Therefore, setting $k=2l$ we have the asymptotic recurrence relation for the even moments:
@@
for large $\nu$. This can potentially be used as an experimental signature of the power-law behavior.
B.3 Wigner quasiprobability distribution
@@
satisfying the canonical commutation relation $\hat{q},\hat{p}]=i$. If the system physically describes a quantum particle trapped in a one-dimensional potential, $\hat{q}$ and $\hat{p}$ are the dimensionless position and momentum operators.
@@
$\begin{split}&\hat{a}\rho\longrightarrow\left(\alpha+\frac{1}{2}\partial_{\alpha^{*}}\right)W(\alpha,\alpha^{*})\equiv D_{1}W(\alpha,\alpha^{*}),\quad\rho\hat{a}^{{\dagger}}\longrightarrow D_{1}^{*}W(\alpha,\alpha^{*})\\ &\rho\hat{a}\longrightarrow\left(\alpha-\frac{1}{2}\partial_{\alpha^{*}}\right)W(\alpha,\alpha^{*})\equiv D_{2}W(\alpha,\alpha^{*}),\quad\hat{a}^{{\dagger}}\rho\longrightarrow D_{2}^{*}W(\alpha,\alpha^{*})\end{split}$ (48)
@@
$0=$ $(\gamma_{1}-\kappa_{1})(1+R\partial_{R})W+\frac{\gamma_{1}+\kappa_{1}}{2}\partial_{R}R\partial_{R}W+(\gamma_{2}-\kappa_{2})\left(4R+2R^{2}\partial_{R}+\frac{1}{2}R\partial_{R}^{2}R\partial_{R}\right)W$ (54)
$+2(\gamma_{2}+\kappa_{2})(R\partial_{R}+R\partial_{R}R\partial_{R})W.$
@@
Appendix C Scaling behavior of the number statistics
Here, we derive analytical predictions for the divergences of the various statistics presented in the main text, such as the mean number of excitation quanta and the second-order correlation function, as we approach the critical point $\delta\to 0^{+}$. While the numerical results shown in the main text are for the special case of $M=2$, the analytical treatment here holds for the general $M$-boson model.
@@ Appendix C Scaling behavior of the number statistics
which vanishes as $\delta^{1-\nu}$ for $0<\nu<1$ and is finite for $\nu>1$. For $k\in\mathbb{Z}^{+}$, the $k$-th moment of the distribution $p(n)$ is given by
$\braket{n^{k}}=A\int_{1}^{1/\delta}dn\,n^{k-\nu}=\frac{A}{k-\nu+1}\left(\frac{1}{\delta^{k-\nu+1}}-1\right)$ (63)
@@
$g^{(2)}(0)=\frac{\braket{n^{2}}-\braket{n}}{\braket{n}^{2}}\sim\begin{cases}\text{finite},\quad&0<\nu<1\\[8.61108pt] \log(1/\delta),\quad&\nu=1\\[8.61108pt] \dfrac{1}{\delta^{\nu-1}},\quad&1<\nu<2\\[8.61108pt] \dfrac{1}{\delta\log^{2}(1/\delta)},\quad&\nu=2\\[8.61108pt] \dfrac{1}{\delta^{3-\nu}},\quad&2<\nu<3\\[8.61108pt] \log(1/\delta),\quad&\nu=3\\[8.61108pt] \text{finite},\quad&\nu>3\end{cases}$ (65)
which is divergent in the regime $1\leq\nu\leq 3$, as $\delta\to 0$. This is consistent with the numerical results shown in the main text.
Appendix D Quantum Liénard system
The quantum Liénard system, proposed in Ref. [19], serves as a quantum mechanical analog to the (classical) Liénard system, which is well-studied in the context of nonlinear dynamics. The general Liénard system is given by the dynamical equations
@@ Appendix D Quantum Liénard system
    $F(q)\equiv\int_{0}^{q}f(u)du$ has exactly one positive root at some $q=a$, negative for $0<q<a$, positive and nondecreasing for $q>a$.
@@
D.1 Quantization of dissipative dynamical systems
Eq. (67) cannot be simulated directly, since the system of equations is not closed in general. Instead, we seek a generator of infinitesimal time translation in the form of a Lindblad master equation. A general construction of the master equation (referred to as quantization) for an arbitrary polynomial $f$ and $g$ can be found in Ref. [19]. For completeness, we briefly summarize the steps involved in the quantization procedure:
@@ D.1 Quantization of dissipative dynamical systems
    Since $\alpha$ is a classical variable (sometimes called a $c$-number), we can change the order of $\alpha$ and $\alpha^{*}$ freely in $f_{1}(\alpha,\alpha^{*})$ and $f_{2}(\alpha,\alpha^{*})$. For convenience, reorder the terms such that all the $\alpha^{*}$ appear on the left of $\alpha$.
@@
    The normal ordering of operators comes from Step 3. The choice of ordering here is simply for convenience. In principle, the master equation can be constructed for any operator ordering.
@@
    for a suitable (and in general non-unique) choice of Hamiltonian $\hat{H}$, jump operators $\hat{c}_{j}$ and decay rates $\Gamma_{j}>0$, such that Eq. (71) is satisfied. The solution to this inverse problem is the subject of Ref. [19], which also proved the existence of a master equation for arbitrary polynomials $f_{1}$ and $f_{2}$.
D.2 Example: Cubic Liénard system
The application of the quantization procedure to the cubic Liénard system (68) is detailed in Ref. [19]. The result is the Lindblad master equation
@@ D.2 Example: Cubic Liénard system
$\hat{H}_{1}=\frac{i}{4}\left[k_{0}-i(k_{1}-1)\right]\hat{a}^{2}-\frac{i}{4}\left(\frac{k_{2}}{2}+ik_{3}\right)\hat{a}^{{\dagger}}\hat{a}^{3}-\frac{i}{16}(k_{2}+ik_{3})\hat{a}^{4}+\text{h.c.},$ (76)
@@
D.3 Additional numerical results
[…]
Figure 4: Steady state populations $\rho_{n,n}$ of the quantum Liénard system against the excitation number $n$. The Hilbert space truncation $N_{T}$ in the simulation ranges from $200$ to $1000$. The black dashed line indicates the power-law fit for $n\in[50,300]$ and $N_{T}=1000$, with the fitted exponent indicated in the figures. The parameters $(k_{0},k_{1},k_{2},k_{3})$ used are: (a) (0.465, 0.353, 1, 0.042), (b) (0.168, 0.821, 1, 0.405), (c) (0.723, 0.372, 1, 0.788).
Next, we study the robustness of the power-law behavior in the quantum Liénard model. The numerical results are shown in Fig. 5. As explained in the main text, we identify the nonlinear dissipation term governed by $k_{2}$ to be the key ingredient for the power law. For example, the power law tail can be observed even when $k_{0}=k_{1}=k_{3}=0$ and $k_{2}=1$ [Fig. 5(a)]. Now, if we artificially set the Hamiltonian $H$ to zero, we get a simplified master equation
$\dot{\rho}=-i[H,\rho]+\frac{1}{2}\mathcal{D}[\hat{a}^{{\dagger}}\hat{a}-\hat{a}^{{\dagger}2}/2]\rho+\frac{3}{8}\mathcal{D}[\hat{a}^{2}]\rho$ (78)
with $H=0$. In this case, the steady state does not exhibit a power law tail [Fig. 5(b)]. We observe that the power law can be recovered by setting $H=-i(\hat{a}^{{\dagger}}\hat{a}^{3}-\hat{a}^{{\dagger}3}\hat{a})/4$ [Fig. 5(c)] or $H=-i(\hat{a}^{4}-\hat{a}^{{\dagger}4})/8$ [Fig. 5(d)]. In other words, this suggests that not all the nonlinear terms in the master equation for the quantum Liénard model are needed for the power-law behavior.
Note that in all the models studied in Fig. 5, the steady state has a two-fold degeneracy corresponding to the $\mathbb{Z}_{2}$ symmetry $\hat{a}\to-\hat{a}$. The results are shown for the steady state in the even parity sector, with similar results observed for the steady state in the odd parity sector.
Figure 5: Steady state populations $\rho_{n,n}$ against the excitation number $n$, in the even parity sector. For odd values of $n$, $\rho_{n,n}=0$ which are omitted from the plots. The Hilbert space truncation $N_{T}$ in the simulation ranges from $200$ to $1000$. The black dashed line indicates the power-law fit for $n\in[50,300]$ and $N_{T}=1000$, with the fitted exponent indicated in the figures. The systems studied are: (a) Quantum Liénard model with $k_{0}=k_{1}=k_{3}=0$ and $k_{2}=1$. (b) Same as (a), but with the Hamiltonian $H=0$. (c) Same as (a), but with the Hamiltonian $H=-i(\hat{a}^{{\dagger}}\hat{a}^{3}-\hat{a}^{{\dagger}3}\hat{a})/4$. (d) Same as (a), but with the Hamiltonian $H=-i(\hat{a}^{4}-\hat{a}^{{\dagger}4})/8$.